Determine the simplest polynomial that has real coefficients and zeroes at x=-2i and x=4 (multiply it out)
Answer by stanbon(75887)

Determine the simplest polynomial that has real coefficients and zeroes at x=-2i and x=4
----
If the coefficients are Real, x = 2i is also a zero.
----
Your Equation:
f(x) = (x-2i)(x+2i)(x-4)
----
= (x^2+4)(x-4)
----
= x^3 -4x^2 + 4x - 16
===========================
Cheers,
Stan H.